# GG-SSMs: Graph-Generating State Space Models

### v1

@@ 2 Related Work
2.1 Limitations of Traditional SSMs
In applications like computer vision and event-based data processing [67, 66], data often exhibits rich multi-dimensional structures with complex spatial relationships that cannot be effectively captured by processing the data as a simple sequence. For instance, pixels in an image or events from an event-based vision sensor have spatial dependencies that span across two or three dimensions. Traditional SSMs struggle to capture these intricate relationships when applied to such data due to their reliance on predetermined sequential paths.
Recent methods like Mamba [19] and VMamba [30] have attempted to address these limitations. However, these approaches still rely on fixed, predetermined paths and lack the adaptability to fully capture the diverse and complex structures present in high-dimensional data. Consequently, they may not efficiently model the intricate dependencies and long-range interactions necessary for tasks that require a comprehensive understanding of spatial relationships. GG-SSMs enable feature propagation across non-sequential paths by constructing graphs that represent the inherent relationships within the data based on feature similarities or dissimilarities.
2.2 Visual State Space Models (VSSMs)
While traditional SSMs excel in capturing sequential dependencies in NLP tasks, they face challenges with the complex spatial structures characteristic of visual data. This limitation restricts their application in visual domains. S4ND [36] pioneered multi-dimensional SSMs by introducing independent 1D SSMs for each dimension. Building on this, Baron et al. [2] extended S4ND to a discrete multi-axial framework, introducing a 2D-SSM spatial layer that effectively generalizes 1D SSMs to 2D settings. Recent advancements focus on various spatial scanning strategies to improve spatial consistency in visual SSMs, including bidirectional [65], four-way [30], continuous [56], zigzag [23], window-based [24], and topology-based scanning [38, 54]. However, these approaches, even if topological, take up a lot of resources and are very slow. We solve all of these problems with our framework.
@@ 3.2.2 State Propagation Along the Graph
The difference between our model and previous models (Mamba [19], VMamba [30]) is shown in Figure 2.
[…]
Figure 2: Comparison of our GG-SSM with Mamba [19] and VMamba [30] models. GG-SSM raw operator processing in (c) is based on Figure 1.

### v2

@@ Abstract
Multimedial Material: For GitHub code, poster and other details visit https://github.com/uzh-rpg/gg_ssms.
@@ 2 Related Work
2.1 Challenges with Sequential-Only SSMs
Data naturally exhibit multi-dimensional structures with complex spatial or spatiotemporal relationships in many computer vision applications and event-based data processing [67, 66]. Traditional State Space Models (SSMs) process signals in a strictly one-dimensional sequence, which makes them ill-suited to capture non-local interactions. For instance, images and event streams include spatial dependencies that span two or more dimensions, but standard SSMs restrict the flow of information along predetermined scanning paths. While these 1D-sequential SSMs excel in tasks such as NLP, their inability to adapt to more complex data topologies leads to poor performance on high-dimensional or multi-variate inputs. Such limitations are particularly pronounced in tasks that demand a global understanding of spatial patterns, such as image classification, optical flow, and dynamic vision sensing.
2.2 Approaches for High-Dimensional Visual SSMs
Researchers have proposed specialized architectures for multi-dimensional data to address the shortcomings of sequential-only processing. S4ND [36] pioneered a multi-dimensional approach by stacking independent 1D SSM modules, enabling limited handling of spatial axes. Baron et al. [2] introduced a discrete multi-axial framework (A2S) incorporating a 2D-SSM spatial layer to better capture spatial dependencies. More recent advances explore sophisticated scanning and correlation techniques to improve spatial consistency, such as bidirectional [65], four-way [30], continuous [56], zigzag [23], window-based [24], and topology-based [38, 54] processing. Among these, Mamba [19] and VMamba [30] propose flexible paths or multi-directional scanning to escape strictly linear reading orders. However, such approaches still rely on fixed or predefined trajectories, limiting their adaptability to the diverse and irregular structures inherent to visual data. Additionally, even approaches that consider specialized topologies tend to become computationally heavy and less scalable.
Our framework, Graph-Generating State Space Models (GG-SSMs), tackles these issues by constructing dynamic graph structures. Rather than imposing fixed scanning paths, GG-SSMs leverage graph edges derived from feature similarities or dissimilarities to capture long-range interactions on the fly. Moreover, an efficient MST-based construction keeps the graph sparse, ensuring scalability and improving the representational power of state propagation. Our experiments show that this improves various vision tasks and time series analyses.
@@ 4.4 Optical Flow Estimation
5 Ablation Study
We conducted an ablation to compare different MST algorithms for graph construction: Chazelle’s MST [6], Kruskal’s, and Prim’s. Table 8 summarizes the results on both ImageNet (Tiny scale) and a widely used time-series benchmark (ETTh2).
Table 8: Ablation of MST algorithms. Accuracy/performance is nearly identical, but Chazelle’s MST consistently offers lower runtime, especially as the dataset size $L$ increases.
Algorithm ImageNet (Tiny) ETTh2 (Horizon=192)
Top-1 (%) Time (s/epoch) MSE Time (s/epoch)
Kruskal 83.5 1.00×1.$0$ 0.356 1.00×1.$0$
Prim 83.4 0.98×0.98$\times$0.98 × 0.358 0.95×0.95$\times$0.95 ×
Chazelle 83.6 0.90×0.90$\times$0.90 × 0.353 0.88×0.$8$
All three algorithms yield comparable model accuracy (within ±$\pm 0.1\%$) but differ in computational efficiency. In particular, Chazelle’s MST runs faster in practice for larger $L$, matching its near-linear theoretical time complexity.
@@ 7 Acknowledgment
[…]
Supplementary Material
8 Chazelle’s MST Algorithm
In this section, we provide a concise overview of Chazelle’s MST algorithm [6] and why it is instrumental to our proposed Graph-Generating State Space Models (GG-SSMs). Although multiple efficient minimum spanning tree (MST) algorithms exist (e.g., Kruskal’s, Prim’s, or Borůvka’s), Chazelle’s algorithm is particularly interesting due to its near-linear time complexity in the general graph setting.
[…]
Figure 2: Chazelle’s MST Overview. Soft heaps allow near-linear sorting of edges. MST edges (in blue) form a spanning structure with no cycles, connecting all vertices using the smallest weights $w$.
8.1 Core Idea and Time Complexity
Chazelle’s MST algorithm belongs to the family of soft heap approaches. Its most prominent feature is achieving a runtime of $\mathcal{O}(E\,\alpha(E,V))$, where
  • $V$ is the number of vertices in the graph,
  • $E$ is the number of edges in the graph,
  • $\alpha(\cdot,\cdot)$ is the inverse Ackermann function, a function that grows extremely slowly (much more slowly than $\log\log n$).
For any practical input size (e.g., up to millions of edges), $\alpha(E,V)$ remains a small constant (typically $\leq 4$). Therefore, the runtime is effectively linear for all real-world purposes.
Algorithmic Outline.
At a high level, Chazelle’s algorithm proceeds by maintaining a specialized priority queue known as a soft heap to handle edge weight comparisons and merges. It selectively corrupts (or perturbs) a small fraction of keys but guarantees a sufficiently accurate ordering to recover the MST. The soft heap structure allows various operations like insertions, extractions, and merges to be executed in approximately constant amortized time, modulo the very slowly growing $\alpha$ factor.
The algorithm can be outlined in three major steps:
  1. 1.
    Sort or partition the edges using the soft-heap structure such that they can be processed in a non-decreasing order of weights.
  2. 2.
    Merge edge sets while extracting edge candidates for the MST. These extractions remain almost linear because each edge is either integrated into the MST structure or discarded.
  3. 3.
    Selective Corruption & Verification: Since the soft heap may slightly perturb edge weights, a verification step ensures that these corrupted weights do not impact the MST correctness. With high probability, only a small fraction of edges require re-checking.
By the end, the edges forming the MST are collected using a Union-Find data structure (or a similar disjoint set data structure), combining efficiency with theoretical guarantees of correctness.
8.2 Practical Implications for GG-SSMs
In our GG-SSM framework, each layer requires constructing an MST on feature embeddings (e.g., pixel or token embeddings) to identify critical connections before performing state propagation. Since the number of edges $E$ can be large in dense graphs, employing Chazelle’s MST algorithm with its near-linear time complexity is particularly appealing:
  • Scalability: For a graph of $L$ vertices (e.g., $L$ embeddings), we can handle $\mathcal{O}(L^{2})$ potential edges in worst-case dense settings or employ faster approximate methods in sparser representations. Either way, Chazelle’s $\mathcal{O}(E\,\alpha(E,V))$ ensures minimal overhead when $E$ is proportional to $L$ or $L\log L$.
  • Uniqueness of MST Paths: MST ensures exactly $L-1$ edges and a unique path between any two nodes. This property is crucial for our GG-SSMs, as it cleanly defines the path by which hidden states propagate. It further limits redundant computations and fosters efficient parameter sharing in state updates.
8.3 Intuition and Benefits
Intuitively, MST-based graph construction identifies the closest (or most similar) neighbors by selecting edges of the smallest weight (lowest dissimilarity). This yields a minimal set of edges connecting all nodes, providing a concise but effective skeleton to diffuse signals across the entire feature set. Consequently, even high-dimensional data with long-range dependencies can be efficiently processed with minimal redundancy. The slow-growing $\alpha(\cdot,\cdot)$ factor ensures that the overhead of constructing and maintaining this structure remains negligible for practical dataset sizes.